# A new spin on polynomial relations among kappa classes

### v1

@@ 2.4.2. Proofs of the main theorems
Finally, for $d=0$ case of Theorem 1.3, we compute the classes explicitly without applying localization, just directly from its definition (see [KL13, KL11] and [GKLS24, Section 1.1]). To this end, we use the following trick: we consider $(\mathbb{P},\mathcal{O}(-1))$ as the normal bundle of the exceptional divisor $\tilde{X}$ in $\tilde{Y}$ obtained as a blow-up at one point of a Calabi-Yau surface $Y$. Then we can just apply the computation of [GP98, Section 2.2]. We have $c_{1}(\tilde{Y})=-[\tilde{X}$], and $\overline{\mathcal{M}}_{g,n}(\tilde{Y},0)]^{\mathrm{vir}}=[\overline{\mathcal{M}}_{g,n}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})$. Then for $p\colon\overline{\mathcal{M}}_{g,n}\times\tilde{Y}\to\overline{\mathcal{M}}_{g,n}$ we have
(56) $)\displaystyle(-1)^{g}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,1}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})\cap ev_{1}^{*}(-[\tilde{X}$$\Big{)}$ $=(-1)^{g}\lambda_{g}\lambda_{g-1};$
(57) $(-1)^{g}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,0}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})\Big{)}$ $=(-1)^{g-1}\lambda_{g}\lambda_{g-2},$
and according to [KL11] the result is still to be twisted by the sign (1)(-1). ∎
@@ 3.1. Topological recursion and its equivariant version
    (61) $dx(\gamma z)=\chi_{\gamma}\,dx(z)\,,\qquad y(\gamma z)=\upsilon_{\gamma}\,y(z)\,,\qquad B(\gamma z_{1},z_{2})=\beta_{\gamma}\,B(z_{1},z_{2})\,$,
@@
An example of a $G$-equivariant spectral curve data is given in Theorem 1.4. It is a $\mathbb{Z}_{2}$-equivariant spectral curve data according to Definition 3.1, up to a minor caveat which does not affect anything. Namely, we have
@@
where $\iota$ denotes the action of the generator of $\mathbb{Z}_{2}$ (that is $\iota(z)=\phi(1,z)$, where $\mathbb{Z}_{2}$ is $\{0,1\}$ with 0 being the group unit); and we have replaced $Q$ with $\epsilon$. For this $\mathbb{Z}_{2}$-equivariant spectral curve $\chi=\upsilon$ is the sign representation and $\beta$ is the trivial representation of $\mathbb{Z}_{2}$. The caveat is that rather than having $y(\iota z)=-y(z)$ as dictated by the definition, we have $y(\iota z)=-y(z)+\mathrm{const}$; but due to the way $y$ enters (59) this extra constant does not affect anything, it gets canceled in the recursion formula. This means that all statements regarding equivariant spectral curves hold for a curve with such a relaxed assumption on $y(\iota(z))$; we will thus slightly extend Definition 3.1 to allow such a relaxed condition, and still call the respective curve an equivariant spectral curve.
@@ 3.2. Expansion of the differentials
(72) $\tilde{\omega}^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{g+n}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}\epsilon^{d}\int_{p_{*}C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{w_{i}^{2k_{i}+2}}dw_{i}.$
@@
(74) $\tilde{\omega}^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{g+n}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}\epsilon^{d}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{x_{i}^{2k_{i}+2}}dx_{i}.$
@@
(75) $\omega^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{n}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}Q^{d}2^{2g-2+n}(-1)^{g+d}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{x_{i}^{2k_{i}+2}}dx_{i}.$
Taking into account the extra sign $(-1)^{g+d}$ that was natural for us to include in the definition of the descendant invariants, we obtain the statement of Theorem 1.4 in the stable range 2g2+n>02g-2+n>0. It extends further to the unstable case $(0,2)$ by a straightforward computation that essentially repeats the proof of Proposition 2.8. Indeed, applying [Jan17, Equation (A.6)] in our case, we get for $(g,n)=(0,2)$

### v2

@@ 2.4.2. Proofs of the main theorems
Finally, for $d=0$ case of Theorem 1.3, we compute the classes explicitly without applying localization, just directly from its definition (see [KL13, KL11] and [GKLS24, Section 1.1]). To this end, we use the following trick: we consider $(\mathbb{P},\mathcal{O}(-1))$ as the normal bundle of the exceptional divisor $\tilde{X}$ in $\tilde{Y}$ obtained as a blow-up at one point of a K3 surface $Y$. Then we can just apply the computation of [GP98, Section 2.2]. We have $c_{1}(\tilde{Y})=-[\tilde{X}$], and $\overline{\mathcal{M}}_{g,n}(\tilde{Y},0)]^{\mathrm{vir}}=[\overline{\mathcal{M}}_{g,n}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})$. Then for $p\colon\overline{\mathcal{M}}_{g,n}\times\tilde{Y}\to\overline{\mathcal{M}}_{g,n}$ we have
(57) $)\displaystyle(-1)^{g+1}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,1}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})\cap ev_{1}^{*}(-[\tilde{X}$$\Big{)}$ $=(-1)^{g-1}\lambda_{g}\lambda_{g-1};$
$(-1)^{g+1}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,0}\times\tilde{Y}]\cap(\lambda_{g}\lambda_{g-1}[\tilde{X}]+\lambda_{g}\lambda_{g-2}[\tilde{X}]^{2})\Big{)}$ $=(-1)^{g}\lambda_{g}\lambda_{g-2}.$
Remark 2.14.
Alternatively, one can use [KL11, Theorem 1.1] in combination with [KL11, Equation (2.6)]. This theorem states that in the case of $d=0$ we have to replace the class $c_{g-1+d}(R^{1}\pi_{*}f^{*}\mathcal{O}(-1))$ in Equation (8) by the class $c_{g-1}(R^{1}\pi_{*}f^{*}\mathcal{O}(-1)-R^{0}\pi_{*}f^{*}\mathcal{O}(1))$. On $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},0)=\overline{\mathcal{M}}_{g,n}\times\mathbb{P}^{1}$ the latter class is equal to $(-1)^{g-1}(\lambda_{g-1}\otimes 1+3\lambda_{g-2}\otimes[\mathrm{pt}])$. Recall that $\overline{\mathcal{M}}_{g,n}(\mathbb{P}^{1},0)]^{\mathrm{vir}}=(-1)^{g-1}(\lambda_{g}\otimes 1-2\lambda_{g-1}\otimes[\mathrm{pt}]) [$GP98, Section 2.1] and in the case $n=1$ we have $ev_{1}^{*}([\mathrm{pt}])=1\otimes[\mathrm{pt}]$. Then for $p\colon\overline{\mathcal{M}}_{g,n}\times\mathbb{P}^{1}\to\overline{\mathcal{M}}_{g,n}$, for $n=0,1$, we have:
(58) $(-1)^{g+1}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,1}\times\mathbb{P}^{1}]\cap(1\otimes[\mathrm{pt}])\cdot(-1)^{g-1}(\lambda_{g}\otimes 1-2\lambda_{g-1}\otimes[\mathrm{pt}])\cdot\quad$
$(-1)^{g-1}(\lambda_{g-1}\otimes 1+3\lambda_{g-2}\otimes[\mathrm{pt}])\Big{)}$ $=(-1)^{g-1}\lambda_{g}\lambda_{g-1};$
$(-1)^{g+1}p_{*}\Big{(}[\overline{\mathcal{M}}_{g,0}\times\mathbb{P}^{1}]\cap(-1)^{g-1}(\lambda_{g}\otimes 1-2\lambda_{g-1}\otimes[\mathrm{pt}])\cdot\quad$
$(-1)^{g-1}(\lambda_{g-1}\otimes 1+3\lambda_{g-2}\otimes[\mathrm{pt}])$ $=(-1)^{g}\lambda_{g}\lambda_{g-2}$
as we have already checked in Equation (57).
@@ 3.1. Topological recursion and its equivariant version
    (62) $dx(\gamma z)=\chi_{\gamma}\,dx(z)\,,\qquad dy(\gamma z)=\upsilon_{\gamma}\,dy(z)\,,\qquad B(\gamma z_{1},z_{2})=\beta_{\gamma}\,B(z_{1},z_{2})\,$,
@@
We also slightly modified (62), namely, we formulate it as a condition on the differential $dy$ comparing to the condition on the function $y$ in [GKL24, Definition 8.1]. This definition is slightly more general, but due to the way $y$ enters (60) all statements regarding equivariant topological recursion hold for a curve with such a relaxed assumption. We need this relaxed condition in the example below.
@@
An example of a $G$-equivariant spectral curve data is given in Theorem 1.4. It is a $\mathbb{Z}_{2}$-equivariant spectral curve data according to Definition 3.1. Namely, we have
@@
where $\iota$ denotes the action of the generator of $\mathbb{Z}_{2}$ (that is $\iota(z)=\phi(1,z)$, where $\mathbb{Z}_{2}$ is $\{0,1\}$ with 0 being the group unit); and we have replaced $Q$ with $\epsilon$. For this $\mathbb{Z}_{2}$-equivariant spectral curve $\chi=\upsilon$ is the sign representation and $\beta$ is the trivial representation of $\mathbb{Z}_{2}$.
@@ 3.2. Expansion of the differentials
(73) $\tilde{\omega}^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{g+d+n+1}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}\epsilon^{d}\int_{p_{*}C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{w_{i}^{2k_{i}+2}}dw_{i}.$
@@
(75) $\tilde{\omega}^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{g+n+1}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}Q^{d}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{x_{i}^{2k_{i}+2}}dx_{i}.$
@@
(76) $\omega^{(g)}_{n}(w_{1},\dots,w_{n})$ $=(-1)^{n}\sum_{k_{1},\dots,k_{n}=0}^{\infty}\sum_{d=0}^{\infty}Q^{d}2^{2g-2+n}(-1)^{g+1}\int_{C_{g,n,d}}\prod_{i=1}^{n}\psi_{i}^{k_{i}}\prod_{i=1}^{n}\frac{(2k_{i}+1)!!}{x_{i}^{2k_{i}+2}}dx_{i}.$
Taking into account the extra sign $(-1)^{g+1}$ that was natural for us to include in the definition of the descendant invariants, we obtain the statement of Theorem 1.4 in the stable range 2g2+n>02g-2+n>0. It extends further to the unstable case $(0,2)$ by a straightforward computation that essentially repeats the proof of Proposition 2.9. Indeed, applying [Jan17, Equation (A.6)] in our case, we get for $(g,n)=(0,2)$